A random sample of 25 tablets of buffered aspirin contains, on average, 325.05 mg

of aspirin per tablet, with a standard deviation of 55/10 mg. Find the 95% tolerance

limits that will contain 90% of the tablet contents for this brand of buffered aspirin.

Assume that the aspirin content is normally distributed.
1
Expert's answer
2021-09-28T14:24:17-0400

We need to construct the "95\%" confidence interval for the population mean "\mu."

The critical value for "\alpha = 0.05" and "df = n-1 = 24" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.063899."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(325.05-2.063899\times\dfrac{5.5}{\sqrt{25}},325.05+2.063899\times\dfrac{5.5}{\sqrt{25}})"

"=(322.780, 327.320)"

Therefore, based on the data provided, the "95\%" confidence interval for the population mean is "322.780<\mu< 327.320," which indicates that we are "95\%" confident that the true population mean "\mu" is contained by the interval "(322.780, 327.320)."